# QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement Learning for Mobile Edge Computing

### v2

@@ V-A Simulation Setup
We consider a MEC environment with 50 MDs and 5 ENs, similar to [12]. We also follow the model presented in [17] to determine the energy consumption. All the parameters are given in Table I. To train the MDs’ neural networks, we adopt a scenario comprising 1000 episodes. Each episode contains 100 time slots, each of length 0.1 second. The QECO algorithm incorporates real-time experience into its training process to continuously enhance the offloading strategy. Specifically, we employ a batch size of 16, maintain a fixed learning rate of 0.001, and set the discount factor $\gamma$ to 0.9. The probability of random exploration gradually decreases from an initial value 1, progressively approaching 0.01, all of which is facilitated by an RMSProp optimizer.
@@
    PGOA [23]: This existing method is a distributed optimization algorithm designed for delay-sensitive tasks in an environment where MDs interact strategically with multiple ENs. We select PGOA as a benchmark method due to its similarity to our work.
Tabelle I: Simulation Parameters
Parameter Value
Computation capacity of MD $f_{i}$ 2.6 GHz
Computation capacity of EN $f_{j}^{\text{E}}$ 42.8 GHz
Transmission capacity of MD $r_{i,j}(t)$ 14 Mbps
Task arrival rate 150 Task/sec
[…]
Deadline of task $\Delta_{i}$ 10 time slots (1 Sec)
[…]
Computation power of EN $p_{j}^{\text{E}}$ 5 W
Transmission power of MD $p_{i}^{\text{T}}$ 2.3 W
Standby power of MD $p_{i}^{\text{I}}$ 0.1 W
V-B Performance Comparison and Convergence
We first evaluate the number of completed tasks when comparing our proposed QECO algorithm with the other four schemes. As illustrated in Fig. 3 (a), the QECO algorithm consistently outperforms the benchmark methods when we vary the task arrival rate. At a lower task arrival rate (i.e., 50), most of the methods demonstrate similar proficiency in completing tasks. However, as the task arrival rate increases, the efficiency of QECO becomes more evident. Specifically, when the task arrival rate increases to 250, our algorithm can increase the number of completed tasks by 73% and 47% compared to RD and PGOA, respectively. Similarly, in Fig. 3 (b), as the number of MDs increases, QECO shows significant improvements in the number of completed tasks compared to other methods, especially when faced with a large number of MDs. When there are 110 MDs, our proposed algorithm can effectively increase the number of completed tasks by at least 34% comparing with other methods. This achievement is attributed to the QECO’s ability to effectively handle unknown workloads and prevent congestion at the ENs.
Figs. 4 (a) and  4 (b) illustrate the overall energy consumption for different values of task arrival rate and the number of MDs, respectively. At the lower task arrival rate, the total energy consumption of all methods is close to each other. The total energy consumption increases when we have a higher task arrival rate. As can be observed from Fig. 4 (a), at task arrival rate 450, QECO effectively reduces overall energy consumption by 18% and 15% compared to RD and PGOA, respectively, as it takes into account the battery level of the MD in its decision-making process. However, it consumes more energy compared to LC and FO because they do not utilize all computing resources. In particular, LC only uses the MD’s computational resources, while FO utilizes the allocated EN computing resources.
@@ V-B Performance Comparison and Convergence
In Fig. 4 (b), an increasing trend in overall energy consumption is observed as the number of MDs increases since the number of resources available in the system increases, which leads to higher energy consumption. The QECO algorithm consistently outperforms RD and PGOA methods in overall energy consumption, especially when there are a large number of MDs. Specifically, QECO demonstrates a 27% and 16% reduction in overall energy consumption compared to RD and PGOA, respectively, when the number of MDs increases to 110.
As shown in Fig. 5 (a), the QECO algorithm maintains a lower average delay compared to other methods as the task arrival rate increases from 50 to 350. Specifically, when the task arrival rate is 200, it reduces the average delay by at least 12% compared to other methods. However, for task arrival rates exceeding 350, QECO may experience a higher average delay compared to some of the other methods. This can be attributed to the fact that the other algorithms drop more tasks while our proposed algorithm is capable of completing a higher number of tasks, potentially leading to an increase in average delay. In Fig. 5 (b), as the number of MDs increases, we observe a rising trend in the average delay. It can be inferred that an increase in computational load in the system can lead to higher queuing delays and computations at ENs. Considering the QECO’s ability to schedule workloads, when the number of MDs increases from 30 to 110, it consistently maintains a lower average delay which is at least 8% less than the other methods.
We further investigate the overall improvement achieved by the QECO algorithm in comparison to other methods in terms of the average QoE. This metric signifies the advantages MDs obtain by utilizing different algorithms. Fig. 6 (a) shows the average QoE for different values of task arrival rate. This figure indicates the superiority of the QECO algorithm in providing MDs with an enhanced experience. Specifically, when the task arrival rate is moderate (i.e., 250), QECO improves the average QoE by 57% and 33% compared to RD and PGOA, respectively. Fig. 6 (b) illustrates the average QoE when we increase the number of MDs. The EN’s workload grows when there are a larger number of MDs, leading to a reduction in the average QoE of all methods except LC. However, QECO effectively manages the uncertain load at the ENs. When the number of MDs increases to 90, QECO achieves at least a 29% higher QoE comparing with the other methods. It is worth noting that although improvements in each of the QoE factors can contribute to enhancing system performance, it is essential to consider the user’s demands in each time slot. Therefore, the key difference between QECO and other methods is that it prioritizes users’ demands, enabling it to strike an appropriate balance among them, ultimately leading to a higher QoE for MDs.
We finally delve into the investigation of the convergence performance of the QECO algorithm, which is shown through the average QoE across episodes in Figs. 7 (a) and 7 (b). We explore the impact of two main hyper-parameters on the convergence speed and the converged result of the proposed algorithm. Fig. 7 (a) illustrates the convergence of the proposed algorithm under different learning rates, where the learning rate regulates the step size per iteration towards minimizing the loss function. The QECO algorithm achieves an average QoE of 0.77 after around 400 episodes when the learning rate is 0.001, indicating relatively rapid convergence. However, with smaller learning rates (e.g., 0.0001) or larger values (e.g., 0.01), a slower convergence is observed. Fig. 7 (b) shows the convergence of the proposed algorithm under different batch sizes, which refer to the number of sampled experiences in each training round. An improvement in convergence performance is observed as the batch size increases from 4 to 16. However, further increasing the batch size from 16 to 32 does not notably enhance the converged QoE or convergence speed. Hence, a batch size of 16 may be more appropriate for training processes.
@@
Abbildung 7: The convergence of the average QoE across episodes under different hyper-parameters: (a) Learning rate; (b) Batch size.
VI Conclusion
In this paper, we focused on addressing the challenge of offloading in MEC systems, where strict task processing deadlines and energy constraints adversely impact system performance. We formulated an optimization problem that aims to maximize the QoE of each MD individually, while QoE reflects the energy consumption and task completion delay. To address the dynamic and uncertain mobile environment, we proposed a QoE-oriented DRL-based computation offloading algorithm called QECO. Our proposed algorithm empowers MDs to make offloading decisions without relying on knowledge about task models or other MDs’ offloading decisions. The QECO algorithm not only adapts to the uncertain dynamics of load levels at ENs, but also effectively manages the ever-changing system environment. Through extensive simulations, we showed that QECO outperforms several established benchmark techniques, while demonstrating a rapid training convergence. Specifically, QECO increases the average user’s QoE by 37% compared to an existing work. This advantage can lead to improvements in key performance metrics, including task completion rate, task delay, and energy consumption, under different system conditions and varying user demands.
@@ VI Conclusion
  • Yousefpour et al. [2019] A. Yousefpour, C. Fung, T. Nguyen, K. Kadiyala, F. Jalali, A. Niakanlahiji, J. Kong, and J. P. Jue, “All one needs to know about fog computing and related edge computing paradigms: A complete survey,” J. Syst. Archit., vol. 98, pp. 289–330, Sep 2019.
@@
  • Shah-Mansouri and Wong [2018] H. Shah-Mansouri and V. W. Wong, “Hierarchical fog-cloud computing for IoT systems: A computation offloading game,” IEEE Internet Things J., vol. 5, no. 4, pp. 3246–3257, May 2018.
@@
  • Wu et al. [2024] L. Wu, P. Sun, H. Chen, Y. Zuo, Y. Zhou, and Y. Yang, “NOMA-enabled multiuser offloading in multicell edge computing networks: A coalition game based approach,” IEEE Trans. Netw. Sci. Eng., vol. 11, no. 2, pp. 2170–2181, Mar 2024.
  • Arulkumaran et al. [2017] K. Arulkumaran, M. P. Deisenroth, M. Brundage, and A. A. Bharath, “Deep reinforcement learning: A brief survey,” IEEE Signal Process Mag, vol. 34, no. 6, pp. 26–38, Nov 2017.
  • Huang et al. [2019] L. Huang, S. Bi, and Y.-J. A. Zhang, “Deep reinforcement learning for online computation offloading in wireless powered mobile-edge computing networks,” IEEE Trans. Mob. Comput., vol. 19, no. 11, pp. 2581–2593, Jul 2019.
@@
  • Zhao et al. [2019] N. Zhao, Y.-C. Liang, D. Niyato, Y. Pei, M. Wu, and Y. Jiang, “Deep reinforcement learning for user association and resource allocation in heterogeneous cellular networks,” IEEE Trans. Wireless Commun., vol. 18, no. 11, pp. 5141–5152, Aug 2019.
  • Tang and Wong [2020] M. Tang and V. W. Wong, “Deep reinforcement learning for task offloading in mobile edge computing systems,” IEEE Trans. Mob. Comput., vol. 21, no. 6, pp. 1985–1997, Nov 2020.
  • Sun et al. [2024] C. Sun, X. Li, C. Wang, Q. He, X. Wang, and V. C. Leung, “Hierarchical deep reinforcement learning for joint service caching and computation offloading in mobile edge-cloud computing,” accepted for publication in IEEE Trans. Services Computing, 2024.
  • Dai et al. [2020] Y. Dai, K. Zhang, S. Maharjan, and Y. Zhang, “Edge intelligence for energy-efficient computation offloading and resource allocation in 5G beyond,” IEEE Trans. Veh. Technol., vol. 69, no. 10, pp. 12 175–12 186, Aug 2020.
  • Huang et al. [2021] H. Huang, Q. Ye, and Y. Zhou, “Deadline-aware task offloading with partially-observable deep reinforcement learning for multi-access edge computing,” IEEE Trans. Netw. Sci. Eng., vol. 9, no. 6, pp. 3870–3885, Sep 2021.
  • Liu et al. [2021] Z. Liu, Y. Zhao, J. Song, C. Qiu, X. Chen, and X. Wang, “Learn to coordinate for computation offloading and resource allocation in edge computing: A rational-based distributed approach,” IEEE Trans. Netw. Sci. Eng., vol. 9, no. 5, pp. 3136–3151, Dec 2021.
  • Gao et al. [2023] Z. Gao, L. Yang, and Y. Dai, “Large-scale computation offloading using a multi-agent reinforcement learning in heterogeneous multi-access edge computing,” IEEE Trans. Mob. Comput., vol. 22, no. 6, pp. 3425–3443, Jan 2023.
  • Gong et al. [2024] Y. Gong, H. Yao, J. Wang, M. Li, and S. Guo, “Edge intelligence-driven joint offloading and resource allocation for future 6G Industrial Internet of Things,” accepted for publication in IEEE Trans. Netw. Sci. Eng., 2024.
  • Liao et al. [2023] L. Liao, Y. Lai, F. Yang, and W. Zeng, “Online computation offloading with double reinforcement learning algorithm in mobile edge computing,” J Parallel Distrib Comput, vol. 171, pp. 28–39, Jan 2023.
  • Mnih et al. [2015] V. Mnih, K. Kavukcuoglu, D. Silver, A. A. Rusu, J. Veness, M. G. Bellemare, A. Graves, M. Riedmiller, A. K. Fidjeland, G. Ostrovski et al., “Human-level control through deep reinforcement learning,” Nature, vol. 518, no. 7540, pp. 529–533, Feb 2015.
  • Hochreiter and Schmidhuber [1997] S. Hochreiter and J. Schmidhuber, “Long short-term memory,” Neural Comput, vol. 9, no. 8, pp. 1735–1780, Sep 1997.
  • Mao et al. [2016] Y. Mao, J. Zhang, and K. B. Letaief, “Dynamic computation offloading for mobile-edge computing with energy harvesting devices,” IEEE J. Sel. Areas Commun., vol. 34, no. 12, pp. 3590–3605, Sep 2016.
@@
  • Wang et al. [2016] Z. Wang, T. Schaul, M. Hessel, H. Hasselt, M. Lanctot, and N. Freitas, “Dueling network architectures for deep reinforcement learning,” in Proc. of International Conference on Machine Learning.   New York, NY, Jun 2016.

### v3

@@ VI-A Simulation Setup
We consider a MEC environment with 50 MDs and 5 ENs, similar to [41]. We also follow the model presented in [20] to determine the energy consumption. All the parameters are given in Table II. Since there is no real dataset due to the challenges of capturing representative samples over extended periods, we use a simulated MEC system that enables DRL agents to continuously gather experience and improve their performance based on system feedback (e.g., QoE). To train the MDs’ neural networks, we adopt a scenario comprising 1000 episodes. Each episode contains 100 time slots, each of length 0.1 second. The QECO algorithm incorporates real-time experience into its training process to continuously enhance the offloading strategy. Specifically, we employ a batch size of 16, maintain a fixed learning rate of 0.001, and set the discount factor $\gamma$ to 0.9. The probability of random exploration gradually decreases from an initial value 1, progressively approaching 0.01, all of which is facilitated by an RMSProp optimizer. The algorithm’s source code is available at [42].
Tabelle II: Simulation Parameters
Parameter Value
Computation capacity of MD $f_{i}$ 2.6 GHz
Computation capacity of EN $f_{j}^{\text{E}}$ 42.8 GHz
Transmission capacity of MD $r_{i,j}(t)$ 14 Mbps
Task arrival rate 150 Task/sec
Size of task $\lambda_{i}(t)$ $\{\text{1.0, 1.1, \ldots, 7.0}\}$ Mbits
Required CPU cycles of task $\rho_{i}(t)$ $\{\text{0.197, 0.297, 0.397}\}$ $\times 10^{3}$
Deadline of task $\Delta_{i}$ 10 time slots (1 Sec)
Percentage of MD $i$’s battery level $\phi_{i}(t)$ $\{\text{25, 50, 75}\}$
Computation power of EN $p_{j}^{\text{E}}$ 5 W
Transmission power of MD $p_{i}^{\text{T}}$ 2.3 W
Standby power of MD $p_{i}^{\text{I}}$ 0.1 W
@@
    PGOA [11]: This existing method is a distributed optimization algorithm designed for delay-sensitive tasks in an environment where MDs interact strategically with multiple ENs.
  5. 5.
    DCDRL [12]: This method is based on the AC framework [43], which underpins many state-of-the-art DRL algorithms, such as DDPG [44], and PPO [45]. DCDRL is designed for distributed computation offloading in a queuing-based MEC environment.
@@
VI-B Performance Comparison and Convergence
We first evaluate the number of completed tasks when comparing our proposed QECO algorithm with the other four schemes. As illustrated in Fig. 3(a), the QECO algorithm consistently outperforms the benchmark methods when we vary the task arrival rate. At a lower task arrival rate (i.e., 50), most of the methods demonstrate similar proficiency in completing tasks. However, as the task arrival rate increases, the efficiency of QECO becomes more evident. Specifically, when the task arrival rate increases to 250, our algorithm can increase the number of completed tasks by 72.1%, 46.8%, and 28.6% compared to RD, PGOA, and DCDRL, respectively. Similarly, in Fig. 3(b), as the number of MDs increases, QECO shows significant improvements in the number of completed tasks compared to other methods, especially when faced with a large number of MDs. When there are 110 MDs, our proposed algorithm can effectively increase the number of completed tasks by at least 22.8% comparing to other methods. This achievement is attributed to the QECO’s ability to effectively handle unknown workloads and prevent congestion at the ENs.
Figs. 4(a) and 4(b) illustrate the overall energy consumption for different values of task arrival rate and the number of MDs, respectively. At the lower task arrival rate, the total energy consumption of all methods is close to each other. The total energy consumption increases when we have a higher task arrival rate. As can be observed from Fig. 4(a), at task arrival rate 450, QECO effectively reduces overall energy consumption by 18.6%, 15.5%, and 13.9% compared to RD, PGOA, and DCDRL, respectively, as it takes into account the energy level of the MD in its decision-making process. However, it consumes more energy compared to LC and FO because they do not utilize all computing resources. In particular, LC only uses the MD’s computational resources, while FO utilizes the allocated EN computing resources. In Fig. 4(b), an increasing trend in overall energy consumption is observed as the number of MDs increases since the number of resources available in the system increases, which leads to higher energy consumption. The QECO algorithm consistently outperforms other methods in overall energy consumption, especially when there are a large number of MDs. Specifically, QECO demonstrates a 27.4%, 16.7%, and 23.5% reduction in overall energy consumption compared to RD, PGOA, and DCDRL, respectively, when the number of MDs increases to 110.
@@ VI-B Performance Comparison and Convergence
As shown in Fig. 5(a), the QECO algorithm maintains a lower average delay compared to other methods as the task arrival rate increases from 50 to 350. Specifically, when the task arrival rate is 200, it reduces the average delay by at least 12.4% compared to other methods. However, for task arrival rates exceeding 350, QECO may experience a higher average delay compared to some of the other methods. This can be attributed to the fact that the other algorithms drop more tasks while our proposed algorithm is capable of completing a higher number of tasks, potentially leading to an increase in average delay. In Fig. 5(b), as the number of MDs increases, we observe a rising trend in the average delay. It can be inferred that an increase in computational load in the system can lead to higher queuing delays and computations at ENs. Considering the QECO’s ability to schedule workloads, when the number of MDs increases from 30 to 110, it consistently maintains a lower average delay which is at least 8.3% less than the other methods.
We further investigate the overall improvement achieved by the QECO algorithm in comparison to other methods in terms of the average QoE. This metric signifies the advantages MDs obtain by utilizing different algorithms. Fig. 6(a) shows the average QoE for different values of the task arrival rate. This figure highlights the superiority of the QECO algorithm in providing MDs with an enhanced experience. At lower task arrival rates (i.e., 50-150), QECO maintains an average QoE of at least 0.72, while the other methods experience a steeper decline, with average QoE dropping to 0–0.56. Specifically, when the task arrival rate is 200, QECO improves the average QoE by at least 65.7% compared to other methods. As task arrival rates increase to 300, the average QoE significantly decreases for all methods due to increased competition for resources in the MEC system. However, QECO still maintains a positive average QoE, while other methods fall to negative values. At higher task arrival rates (i.e., 350–450), QECO achieves 42.6%, 28.5%, and 22.6% improvement in QoE compared to RD, PGOA, and DCDRL, respectively.
Fig. 6(b) illustrates the average QoE when we increase the number of MDs. The EN’s workload grows when there are a larger number of MDs, leading to a reduction in the average QoE of all methods except LC. However, QECO effectively manages the uncertain load at the ENs. When the number of MDs increases from 30 to 110, QECO consistently maintains at least a 24.8% higher QoE compared to the other methods. Specifically, at a moderate number of 70 MDs, QECO achieves an average QoE of 0.57, showing a 70.3% and 44.7% improvement compared to PGOA and DCDRL, respectively. It is worth noting that although improvements in each of the QoE factors can contribute to enhancing system performance, it is essential to consider the user’s demands in each time slot. Therefore, the key difference between QECO and other methods is that it prioritizes users’ demands, enabling it to strike an appropriate balance among them, ultimately leading to a higher QoE for MDs.
@@
Abbildung 7: The convergence of the average QoE across episodes under different DQN-based methods.
@@
Abbildung 8: The convergence of the average QoE across episodes under different hyper-parameters: (a) Learning rate; (b) Batch size.
We finally delve into the investigation of the convergence performance of the QECO algorithm in Fig. 7 and Fig. 8. To validate the effectiveness of the QECO algorithm, we assess its convergence rate compared to the vanilla DQN and DDQN configurations [8], measured by the average QoE across episodes. As shown in Fig. 7, the MD’s network progressively learns efficient policies over the episodes, ultimately stabilizing as it approaches convergence. Specifically, DQN, DDQN, and our proposed QECO algorithm converge after approximately 650, 550, and 450 iterations, respectively. The QECO algorithm demonstrates the faster convergence while achieving a higher average QoE than the other methods. This underscores the beneficial impact of workload prediction by the LSTM network and highlights its effectiveness in efficiently utilizing the computing resources of MDs and ENs.
Furthermore, we explore the impact of two main hyper-parameters on the convergence speed and the converged result of the proposed algorithm. Fig. 8(a) illustrates the convergence of the proposed algorithm under different learning rates, where the learning rate regulates the step size per iteration towards minimizing the loss function. The QECO algorithm achieves an average QoE of 0.75 when the learning rate is 0.001, indicating relatively rapid convergence. However, with smaller learning rates (e.g., 0.0001) or larger values (e.g., 0.01), a slower convergence is observed. Fig. 8(b) shows the convergence of the proposed algorithm under different batch sizes, which refer to the number of sampled experiences in each training round. An improvement in convergence performance is observed as the batch size increases from 4 to 16. However, further increasing the batch size from 16 to 32 does not notably enhance the converged QoE or convergence speed. Hence, a batch size of 16 may be more appropriate for training processes.
VII Conclusion
In this paper, we focused on addressing the challenge of offloading in MEC systems, where strict task processing deadlines and energy constraints adversely impact system performance. We formulated an optimization problem that aims to maximize the QoE of each MD individually, while QoE reflects the energy consumption and task completion delay. To address the dynamic and uncertain mobile environment, we proposed a QoE-oriented DRL-based computation offloading algorithm called QECO. Our proposed algorithm empowers MDs to make offloading decisions without relying on knowledge about task models or other MDs’ offloading decisions. The QECO algorithm not only adapts to the uncertain dynamics of load levels at ENs, but also effectively manages the ever-changing system environment. Through extensive simulations, we showed that QECO outperforms several established benchmark techniques, while demonstrating a rapid training convergence. Specifically, QECO increases the average user’s QoE by 29.8% – 37.1% compared to several existing algorithms. This advantage can lead to improvements in key performance metrics, including task completion rate, task delay, and energy consumption, under different system conditions and varying user demands.
@@ VII Conclusion
  • Shah-Mansouri and Wong [2018] H. Shah-Mansouri and V. W. Wong, “Hierarchical fog-cloud computing for IoT systems: A computation offloading game,” IEEE Internet Things J., vol. 5, no. 4, pp. 3246–3257, Aug 2018.
@@
  • Wu et al. [2024a] L. Wu, P. Sun, H. Chen, Y. Zuo, Y. Zhou, and Y. Yang, “NOMA-enabled multiuser offloading in multicell edge computing networks: A coalition game based approach,” IEEE Trans. Netw. Sci. Eng., vol. 11, no. 2, pp. 2170–2181, Mar 2024.
  • Mnih et al. [2015] V. Mnih et al., “Human-level control through deep reinforcement learning,” Nature, vol. 518, no. 7540, pp. 529–533, Feb. 2015.
  • Hasselt et al. [2016] H. V. Hasselt, A. Guez, and D. Silver, “Deep reinforcement learning with double Q-learning,” in Proc. AAAI Conf. Artif. Intell., Phoenix, AZ, USA, Feb. 2016, pp. 2094–2100.
  • Wang et al. [2016] Z. Wang, T. Schaul, M. Hessel, H. V. Hasselt, M. Lanctot, and N. de Freitas, “Dueling network architectures for deep reinforcement learning,” in Proc. Int. Conf. Mach. Learn., New York, NY, USA, Jun. 2016, pp. 1995–2003.
  • Hochreiter and Schmidhuber [1997] S. Hochreiter and J. Schmidhuber, “Long short-term memory,” Neural Comput., vol. 9, no. 8, pp. 1735–1780, Sep. 1997.
  • Qiu et al. [2021] X. Qiu, W. Zhang, W. Chen, and Z. Zheng, “Distributed and collective deep reinforcement learning for computation offloading: A practical perspective,” IEEE Trans. Parallel Distrib. Syst., vol. 32, no. 5, pp. 1085–1101, May 2021.
  • Zhang et al. [2024] B. Zhang, F. Xiao, and L. Wu, “Offline reinforcement learning for asynchronous task offloading in mobile edge computing,” IEEE Trans. Netw. Serv. Manag., vol. 21, no. 1, pp. 939–952, Feb 2024.
  • Sun et al. [2024] C. Sun, X. Li, C. Wang, Q. He, X. Wang, and V. C. Leung, “Hierarchical deep reinforcement learning for joint service caching and computation offloading in mobile edge-cloud computing,” accepted for publication in IEEE Trans. Services Computing, Aug 2024.
  • Liu et al. [2022a] Y. Liu, J. Yan, and X. Zhao, “Deep reinforcement learning based latency minimization for mobile edge computing with virtualization in maritime uav communication network,” IEEE Trans. Veh. Technol., vol. 71, no. 4, pp. 4225–4236, Apr 2022.
  • Tang et al. [2023] H. Tang, H. Wu, G. Qu, and R. Li, “Double deep Q-network based dynamic framing offloading in vehicular edge computing,” IEEE Trans. Netw. Sci. Eng., vol. 10, no. 3, pp. 1297–1310, Jun 2023.
  • Li et al. [2022] X. Li, L. Huang, H. Wang, S. Bi, and Y.-J. A. Zhang, “An integrated optimization-learning framework for online combinatorial computation offloading in mec networks,” IEEE Wirel. Commun., vol. 29, no. 1, pp. 170–177, Feb 2022.
  • Wei et al. [2024] Z. Wei, B. Li, R. Zhang, X. Cheng, and L. Yang, “Many-to-many task offloading in vehicular fog computing: A multi-agent deep reinforcement learning approach,” IEEE Trans. Mob. Comput., vol. 23, no. 3, pp. 2107–2122, Mar 2024.
  • Munir et al. [2021] M. S. Munir, N. H. Tran, W. Saad, and C. S. Hong, “Multi-agent meta-reinforcement learning for self-powered and sustainable edge computing systems,” IEEE Trans. Netw. Serv. Manag., vol. 18, no. 3, pp. 3353–3374, Sep 2021.
  • Dai et al. [2020] Y. Dai, K. Zhang, S. Maharjan, and Y. Zhang, “Edge intelligence for energy-efficient computation offloading and resource allocation in 5G beyond,” IEEE Trans. Veh. Technol., vol. 69, no. 10, pp. 12 175–12 186, Oct 2020.
  • Chouikhi et al. [2024] S. Chouikhi, M. Esseghir, and L. Merghem-Boulahia, “Energy-efficient computation offloading based on multi-agent deep reinforcement learning for Industrial Internet of Things systems,” IEEE Internet Things J., vol. 11, no. 7, pp. 12 228–12 239, Apr 2024.
  • Wu et al. [2024b] G. Wu, X. Chen, Y. Shen, Z. Xu, H. Zhang, S. Shen, and S. Yu, “Combining Lyapunov optimization with actor-critic networks for privacy-aware IIoT computation offloading,” IEEE Internet Things J., vol. 11, no. 10, pp. 17 437–17 452, May 2024.
  • Huang et al. [2022] H. Huang, Q. Ye, and Y. Zhou, “Deadline-aware task offloading with partially-observable deep reinforcement learning for multi-access edge computing,” IEEE Trans. Netw. Sci. Eng., vol. 9, no. 6, pp. 3870–3885, Nov 2022.
  • Liao et al. [2023] L. Liao, Y. Lai, F. Yang, and W. Zeng, “Online computation offloading with double reinforcement learning algorithm in mobile edge computing,” J. Parallel Distrib. Comput., vol. 171, pp. 28–39, Jan 2023.
  • Wu et al. [2023a] G. Wu, Z. Xu, H. Zhang, S. Shen, and S. Yu, “Multi-agent DRL for joint completion delay and energy consumption with queuing theory in mec-based IIoT,” J. Parallel Distrib. Comput., vol. 176, pp. 80–94, Jun 2023.
  • Gong et al. [2024] Y. Gong, H. Yao, J. Wang, M. Li, and S. Guo, “Edge intelligence-driven joint offloading and resource allocation for future 6G Industrial Internet of Things,” accepted for publication in IEEE Trans. Netw. Sci. Eng., Dec 2024.
  • Liu et al. [2022b] Z. Liu, Y. Zhao, J. Song, C. Qiu, X. Chen, and X. Wang, “Learn to coordinate for computation offloading and resource allocation in edge computing: A rational-based distributed approach,” IEEE Trans. Netw. Sci. Eng., vol. 9, no. 5, pp. 3136–3151, Sep 2022.
  • She et al. [2024] H. She, L. Yan, and Y. Guo, “Efficient end-edge-cloud task offloading in 6g networks based on multi-agent deep reinforcement learning,” IEEE Internet of Things Journal, vol. 11, no. 11, pp. 20 260–20 270, Jun 2024.
  • Gao et al. [2023] Z. Gao, L. Yang, and Y. Dai, “Large-scale computation offloading using a multi-agent reinforcement learning in heterogeneous multi-access edge computing,” IEEE Trans. Mob. Comput., vol. 22, no. 6, pp. 3425–3443, Jun 2023.
  • Wu et al. [2024c] G. Wu, X. Chen, Z. Gao, H. Zhang, S. Yu, and S. Shen, “Privacy-preserving offloading scheme in multi-access mobile edge computing based on MADRL,” J. Parallel Distrib. Comput., vol. 183, p. 104775, Jan 2024.
@@
  • Wu et al. [2023b] G. Wu, H. Wang, H. Zhang, Y. Zhao, S. Yu, and S. Shen, “Computation offloading method using stochastic games for software-defined-network-based multiagent mobile edge computing,” IEEE Internet Things J., vol. 10, no. 20, pp. 17 620–17 634, Oct 2023.
  • Chen et al. [2021] J. Chen, H. Xing, Z. Xiao, L. Xu, and T. Tao, “A DRL agent for jointly optimizing computation offloading and resource allocation in MEC,” IEEE Internet Things J, vol. 8, no. 24, pp. 17 508–17 524, Dec 2021.
  • Zhao et al. [2019] N. Zhao, Y.-C. Liang, D. Niyato, Y. Pei, M. Wu, and Y. Jiang, “Deep reinforcement learning for user association and resource allocation in heterogeneous cellular networks,” IEEE Trans. Wireless Commun., vol. 18, no. 11, pp. 5141–5152, Nov 2019.
  • Wang et al. [2020] F. Wang, F. Wang, J. Liu, R. Shea, and L. Sun, “Intelligent video caching at network edge: A multi-agent deep reinforcement learning approach,” in Proc. IEEE INFOCOM, Toronto, Canada, Aug. 2020, pp. 2499–2508.
  • Zhang et al. [2019] C. Zhang et al., “Toward edge-assisted video content intelligent caching with long short-term memory learning,” IEEE Access, vol. 7, pp. 152 832–152 846, 2019.
  • Mao et al. [2016] Y. Mao, J. Zhang, and K. B. Letaief, “Dynamic computation offloading for mobile-edge computing with energy harvesting devices,” IEEE J. Sel. Areas Commun., vol. 34, no. 12, pp. 3590–3605, Dec 2016.
@@
  • Lu et al. [2020] H. Lu, X. He, M. Du, X. Ruan, Y. Sun, and K. Wang, “Edge QoE: Computation offloading with deep reinforcement learning for Internet of Things,” IEEE Internet Things J, vol. 7, no. 10, pp. 9255–9265, Oct 2020.
  • Tang and Wong [2022] M. Tang and V. W. Wong, “Deep reinforcement learning for task offloading in mobile edge computing systems,” IEEE Trans. Mob. Comput., vol. 21, no. 6, pp. 1985–1997, Jun 2022.
  • Rahmaty [2024] I. Rahmaty, “QECO source code GitHub repository,” [Online]. Available: https://github.com/ImanRHT/QECO, 2024.
  • Konda and Tsitsiklis [1999] V. Konda and J. Tsitsiklis, “Actor-critic algorithms,” in Proc. Adv. Neural Inf. Process. Syst., vol. 12, apr 1999, pp. 1008–1014.
  • Lillicrap et al. [2016] T. P. Lillicrap et al., “Continuous control with deep reinforcement learning,” in Proc. ICLR, San Juan, Puerto Rico, may 2016, pp. 1–10.
  • Schulman et al. [2017] J. Schulman, F. Wolski, P. Dhariwal, A. Radford, and O. Klimov, “Proximal policy optimization algorithms,” arXiv:1707.06347, jul 2017.
[Uncaptioned image] Iman Rahmaty received the M.Sc. degree in computer science and engineering from Sharif University of Technology, Tehran, Iran, in 2022, he was a Research Assistant with the Performance and Dependability Laboratory. He was with the Department of Electrical Engineering, Sharif University of Technology, in 2023, where he is currently a Research Engineer in the EdgeAI laboratory. His research interests include applying AI/ML-driven solutions, including reinforcement learning, federated learning, and meta-learning to address the challenges and complexities of distributed systems, Internet of Things, mobile edge computing, and wireless networks.
[Uncaptioned image] Hamed Shah-Mansouri (Member, IEEE) received the B.Sc., M.Sc., and Ph.D. degrees in electrical engineering from the Sharif University of Technology, Tehran, Iran, in 2005, 2007, and 2012, respectively. He was a Postdoctoral Research and a Teaching Fellow with the University of British Columbia, Vancouver, BC, Canada. He was with the Department of Electrical Engineering, Sharif University of Technology, in 2019 where he is currently an Assistant Professor. He has served as the publication co-chair for the IEEE Canadian Conference on Electrical and Computer Engineering 2016 and IEEE International Conference on Systems, Man and Cybernetics 2020 and as the technical program committee (TPC) member for several conferences including the IEEE Global Communications Conference (GLOBECOM) 2015 and the IEEE Vehicular Technology Conference (VTC–Fall) 2016, 2017, and 2018. His research interests include the application of optimization and machine learning in computer networks, and Internet of Things.
[Uncaptioned image] Ali Movaghar (Life Senior Member, IEEE) received the M.S. and Ph.D. degrees in computer, information, and control engineering from the University of Michigan, Ann Arbor, MI, USA, in 1979 and 1985, respectively. He is an Emeritus Professor in the Department of Computer Engineering at Sharif University of Technology, Tehran, Iran, where he has been a faculty member since 1993. Currently, he is a Visiting Professor at the University of Michigan, Ann Arbor, and has previously held visiting positions at the Institut National de Recherche en Informatique et en Automatique (INRIA), Paris, France, in 1984, and in the Department of Electrical Engineering and Computer Science at the University of California, Irvine, in 2011. He also worked with AT&T Information Systems in Naperville, Illinois in 1985-1986, and taught with the University of Michigan, Ann Arbor, in 1987-1989. He is a senior member of the ACM. His research interests include performance and dependability modeling, formal verification of wireless networks, distributed real-time systems, and Internet of Things.